A Bibliography of Papers in Lecture Notes in Computer Science (2004) (Part 1 of 4)

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

04 February 2014 Version 1.05

Title word cross-reference

3 [53, 24]. 5 [43]. μoνo2πoλυ [20]. Ω [26, 18].

-Round [53].

0 [54].

1 [35].

2.0 [3].

Abstract [38]. Adaptive [80, 87]. Adaptively [56]. Advanced [36]. Against [59]. Algebra [7]. Algebraic [60]. Algorithms [8]. Alphabet [23]. Alternative [81, 7]. Among [88]. Analysis [11, 15]. Analysis-Based [15]. Ancient [22]. Animation [30]. Animations [30]. Anonymous [40]. Application [55, 78, 94]. Applications [56, 50, 81, 79, 89, 91]. Approach [92]. Approaches [14, 5]. Approximations [37]. Arabic [28, 27]. Architecture [73]. Architectures [81]. Archive [29]. Arithmetic [2]. Assumptions [53]. Asymptotically [46]. Attack [61, 59]. Attacks [52, 50, 60]. Authenticated [65]. Automated [34]. Available [85]. Aware [73, 92]. Awareness [91].

Backmatter [16]. Bare [51]. Based [96, 63, 65, 67, 84, 75, 20, 94, 15, 83, 46]. Basque [19]. Beyond [32, 36]. Bi [38]. Bi-linear [38]. Bilinear [40, 41]. Binary [44]. Bluetooth [61]. Book [16, 1]. Bounded [48, 6, 64]. Bounded-Error [6]. Broadcast [70]. Bulletin [29].

C [3]. C-XSC [3]. Cascade [66]. Cascaded [55]. Case [29, 4, 19]. CBC [66]. Ciphers [60]. Ciphertexts [47]. CJK [26]. Classification [41]. Clocked [59]. COCONUT [13]. Code [34]. Coins [42]. Collaborative [89, 71]. Collisions [54, 42]. Commitments [52, 50]. Communication [46]. Communications [58]. Comparing [5]. Comparison [35]. Complete [41]. Complexity [60]. Components [94]. Composable [56]. Composition [77, 86]. Compress [47]. Compressed [45]. Computation [9, 70, 57, 7]. Computational [12]. Computations [4, 15]. Computer [7]. Computing [3, 49, 91]. Concurrent [51, 50]. Constant [51]. Constant-Round [51]. Constructions [55]. Content [72, 84]. Content-Based [84]. Context [6]. Conversion [20]. Cooperative [6]. Coordinated [86]. Core [41]. Correlated [70]. Correlation [61, 59]. Creating [35]. Credentials [40]. Cryptanalysis [38]. Cryptography [56, 46].

D [24]. Data [80, 89]. Databases [68]. Datamining [68]. Derivation [66]. Deseret [23]. Deterministic [49]. Developing [94]. Devices [67, 89]. Dictionary [21]. Different [5]. Digital [17]. Digits [48]. Directed [76]. Directly [30]. Discrete [48]. Dissemination [82]. Distance [9]. Distributed [95]. Do [42]. Documents [28, 17]. Dynamic [27, 93]. Dynamically [90].

e-Documents [28]. E0 [61]. Edge [81]. Edge-Server [81]. Editing [25]. Efficient [67, 95]. Electronic [58]. Embedding [30]. Encryption [63, 62]. Engine [17]. English [21]. Enterprise [81]. Environment [31, 87]. Equivalent [49]. Error [6]. Esthetics [34]. Estimation [6]. Evaluating [72]. Evaluation [75]. Evolving [88]. Exchange [65]. Experimental [75]. Exploration [71]. Exploring [77]. Exponent [53]. Extended [38, 3]. Extending [93]. Extraction [66].

Factoring [49]. Fast [60]. Faster [61]. Features [36]. Feistel [38, 43]. Feldman [56]. Fields [48]. Filtered [59]. Finding [42]. Finite [9, 48]. Flexible [93, 17]. Floating [15]. Floating-Point [15]. Flood [76]. Flood-Routing [76]. Fonts [27, 35, 36]. Foreseer [73]. Four [4]. Framework [76]. Frontmatter [1]. Functions [41, 42, 55].

Game [92]. Ganymed [79]. Generalizing [19]. Generated [30]. Generator [61]. Generators [59]. Geodesy [12]. Geometry [9]. Global [13]. Glyphs [36]. Gossip [75]. Gossip-Based [75]. Greek [22, 20]. Group [39]. Groups [67]. Guaranteed [7]. Guiding [74]. GUST [29].

Hard [41]. Hard-Core [41]. Hash [42, 55]. Heterogeneous [89]. Highly [85]. Highly-Available [85]. HMAC [66]. Hybrid [62]. Hyphenation [22].

Identifiability [7]. Identity [63].

Implementations [75, 78]. Improved [59]. Independent [96]. Independently [88]. InfoBeacons [74]. Information [74, 82, 71]. Infrastructure [78]. Integrated [31]. Interactive [4, 25, 64]. Interoperability [88]. Intersection [9]. Interval [10, 8, 5, 2]. iOverlay [78]. IPAKE [65]. Irregular [59]. Isomorphisms [65]. Iterated [55]. iTEXMac [31].

J2EE [93]. Java [20, 95]. Java-Based [20].

JavaBeans [81].

Key [65, 51, 66, 49]. Keystream [61, 59]. Keyword [73]. Knowledge [53, 51, 52, 50]. Knowledge-of-Exponent [53].

Language [19, 2]. Languages [26]. Liberty [58]. Libraries [4, 2]. Library [3]. Lightweight [78]. Linear [37, 38]. Locality [73]. Locality-Aware [73]. Location [91]. Logarithm [48]. Low [67]. Low-State [67].

Mac [31]. Man [50]. Man-in-the-Middle [50]. Management [93]. Managing [94, 33]. Maps [40, 33]. Marāth¯ ı [21]. Marāth¯ ı-English [21]. Markup [25]. Mars [71]. Mathematical [28, 27]. MathML [25]. Meghdoot [84]. Message [69]. Method [30]. Middle [50]. Middleware [78, 80, 92, 89, 91, 90, 71, 85]. MiddleWhere [91]. Migrating [29]. Millennium [17]. Mission [71]. MlBibTEX [32]. Model [96, 72, 51, 64]. Modeling [8]. Models [7]. Modern [22]. Modes [66]. Monotonic [20]. Moving [18]. Multi [70, 50]. Multi-party [70]. Multi-trapdoor [50]. Multibody [8]. Multicollisions [55]. Multiple [37, 5].

NASA [71]. Near [54]. Near-Collisions [54]. Need [42]. Networks [84, 76]. Non [64]. Non-interactive [64]. Nonlinear [6]. Novel [73, 14]. Numerical [11, 14, 7].

Object [18]. Object-Oriented [18]. Ontology [94, 83]. Ontology-Based [94, 83]. OOP [2]. Optimal [57, 69, 46]. Optimization [13]. Optimizations [77]. Oracles [63]. Oriented [18]. Overlay [78].

P2P [84]. Packages [24, 5]. Pairings [45]. Paradigm [62]. Parameter [6]. Partitioned [68]. Party [57, 70]. Password [65]. Password-Based [65]. Patterns [22].

PDF [30]. pdfTEX-Generated [30]. Peer [73, 72, 75]. Peer-to-Peer [73, 72]. Subscribe [84, 83, 85]. Perfectly [69]. Performance [11, 77]. Platform [96, 18]. Point [15]. Polynomial [49]. Polytonic [20]. Portable [95]. Portal [71]. Power [92]. Precision [9, 5]. Preserving [68]. Privacy [68]. Privacy-Preserving [68]. Problem [48]. Problems [12]. Programmable [90]. Project [13]. Proofs [52, 50]. Propagation [85]. Protocols [53]. Pseudo [70]. Pseudo-signatures [70]. Public [51, 42]. Public-Key [51]. Publish [84, 83, 85]. Publish/Subscribe [84, 83, 85].

Quantum [52]. Queries [74].

Rabin [47]. Random [63, 43]. Randomness [66, 70]. Reconfigurable [90]. Reliable [9]. Replication [80, 79]. Representations [44]. Resettable [51]. Resource [93, 87]. Resources [33]. Result [12, 14]. Results [35]. Review [24]. Revisited [44]. Revocation [67]. Rewriting [60]. Road [42]. Round [53, 51, 57]. Round-Optimal [57]. Rounds [43]. Routing [76]. Rovers [71]. RSA [49].

SäferTEX [34]. Sampling [75]. Scalable [79]. Scheme [62]. Schemes [40, 38, 43]. Scientific [3]. Searches [73]. Searching [72]. Secret [42, 49]. Secure [56, 63, 50, 42, 57, 69]. Security [58, 43]. Sensor [76]. Server [93, 81, 94]. Service [77, 75]. Services [87, 88, 90, 71, 86]. SHA [54]. SHA-0 [54]. Sharing [87]. Short [39]. Signature [40]. Signatures [39, 47, 70]. Signed [44]. Singular [10]. Small [89]. Software [14, 94]. Soundness [51]. Source [34]. Sources [74, 35]. State [67, 6]. Static [15]. Storage [64]. Stores [89]. Stream [60]. String [52]. Studies [4]. Study [19]. Subscription [85]. Sum [48]. Sum-of-Digits [48]. Support [19, 2].

Switching [11]. SyD [89]. Syntax [25]. System [73, 83]. Systems [10, 8, 11, 4, 6].

Techniques [72, 35]. Telecommunication [11]. Testbed [89]. Testing [7]. Their [50]. Theoretic [92]. Threads [95]. Threshold [56]. Time [49]. Timestamping [64]. Tools [4, 35]. Topic [33]. Torus [46]. Torus-Based [46]. Transactional [79]. Transformation [96]. Transmission [69]. Transparent [82]. trapdoor [50]. Tree [67]. Tree-Based [67]. TRIPLE [96]. Two [57]. Two-Party [57]. Type [35]. Typeset [21]. Typesetting [23, 26, 34]. Typographic [36]. Typography [17].

Ubiquitous [91]. Universally [56]. Universally-Composable [56]. Unstructured [75]. Using [21, 9, 66, 25].

Validation [15]. Variables [60]. Verification [12, 14]. Verified [11, 4]. Vertically [68]. VSS [56].

Web [77, 87, 79, 88, 71, 86]. Wireless [76]. Without [63]. Withstanding [52].

X [31]. XML [29, 33]. XSC [3].

Zero [53, 51, 52]. Zero-Knowledge [53, 52].

References

Anonymous:2004:BF

[1] Anonymous. Book frontmatter. Lecture Notes in Computer Science, 2991: ??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://www.springerlink.com/
media/public/issuefrontmatter/9/3/
9349g6jh2l69.pdf.

vonGudenberg:2004:OIA

[2] Jürgen Wolff von Gudenberg. OOP and interval arithmetic — language support and libraries. Lecture Notes in Computer Science, 2991:1–14, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

springeronline.com/3-540-21260-4.

Hofschuster:2004:CXC

[3] Werner Hofschuster and Walter Krämer. C-XSC 2.0 — A C++ library for extended scientific computing. Lecture Notes in Computer Science, 2991:15–35, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

http://www.springeronline.com/3-
540-21260-4.

Kearfott:2004:LTI

[4] R. Baker Kearfott, Markus Neher, Shin’ichi Oishi, and Fabien Rico. Libraries, tools, and interactive systems for verified computations four case studies. Lecture Notes in Computer Science, 2991:36–63, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

springeronline.com/3-540-21260-4.

Grimmer:2004:MPI

[5] Markus Grimmer, Knut Petras, and Nathalie Revol. Multiple precision interval packages: Comparing different approaches. Lecture Notes in Computer Science, 2991:64–90, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

springeronline.com/3-540-21260-4.

Kieffer:2004:NPS

[6] Michel Kieffer and Eric Walter. Nonlinear parameter and state estimation for cooperative systems in a boundederror context. Lecture Notes in Computer Science, 2991:107–123, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

springeronline.com/3-540-21260-4.

Walter:2004:GNC

[7] Eric Walter, Isabelle Braems, Luc Jaulin, and Michel Kieffer. Guaranteed numerical computation as an alternative to computer algebra for testing models for identifiability. Lecture Notes in Computer Science, 2991:124–131, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

springeronline.com/3-540-21260-4.

Auer:2004:IAM

[8] Ekaterina Auer, Andrés Kecskeméthy, Martin Tändl, and Holger Traczinski. Interval algorithms in modeling of multibody systems. Lecture Notes in Computer Science, 2991:132–159, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

springeronline.com/3-540-21260-4.

Buhler:2004:RDI

[9] Katja Bühler, Eva Dyllong, and Wolfram Luther. Reliable distance and intersection computation using finite precision geometry. Lecture Notes in Computer Science, 2991:160–190, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

springeronline.com/3-540-21260-4.

Alefeld:2004:SIS

[10] Götz Alefeld and Günter Mayer. On singular interval systems. Lecture Notes in Computer Science, 2991:191–197, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://www.springeronline.com/3-
                                        540-21260-4.

Fausten:2004:VNA

[11] Daniela Fausten and Gerhard Haßlinger. Verified numerical analysis of the performance of switching systems in telecommunication. Lecture Notes in Computer Science, 2991:206–225, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

                                        springeronline.com/3-540-21260-4.

Borovac:2004:RVC

[12] Stefan Borovac and Gerhard Heindl. Result verification for computational problems in geodesy. Lecture Notes in Computer Science, 2991:226–242, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://www.springeronline.com/3-
                                        540-21260-4.

Schichl:2004:GOC

[13] Hermann Schichl. Global optimization in the COCONUT Project. Lecture Notes in Computer Science, 2991:243– 249, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL

                                        http://www.springeronline.com/3-
                                        540-21260-4.

Granvilliers:2004:NAN

[14] Laurent Granvilliers, Vladik Kreinovich, and Norbert Müller. Novel approaches to numerical software with result verification. Lecture Notes in Computer Science, 2991:274–305, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

                                        springeronline.com/3-540-21260-4.

Putot:2004:SAB

[15] Sylvie Putot, Eric Goubault, and Matthieu Martel. Static analysis-based validation of floating-point computations. Lecture Notes in Computer Science, 2991:306–313, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://www.

springeronline.com/3-540-21260-4.

Anonymous:2004:BB

[16] Anonymous. Book backmatter. Lecture Notes in Computer Science, 2991: ??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://www.springerlink.com/
media/public/issuebackmatter/9/3/
9349g6jh2l69.pdf.

Loverdos:2004:DTN

[17] Christos K. K. Loverdos and Apostolos Syropoulos. Digital typography in the new millennium: Flexible documents by a flexible engine. Lecture Notes in Computer Science, 3130:1–16, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plaice:2004:MOO

[18] John Plaice, Yannis Haralambous, Paul Swoboda, and Gábor Bella. Moving Ω to an object-oriented platform. Lecture Notes in Computer Science, 3130:17–26, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Perez:2004:BCS

[19] Jagoba Arias Pérez, Jesús Lázaro, and Juan M. Aguirregabiria. Basque: A case study in generalizing language support. Lecture Notes in Computer Science, 3130: 27–33, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Likos:2004:JBC

[20] Johannis Likos. μoνo2πoλυ: Javabased conversion of monotonic to polytonic Greek. Lecture Notes in Computer Science, 3130:34–54, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Athale:2004:ULT

[21] Manasi Athale and Rahul Athale. Using L^{A}TEX to typeset a Marāt.h¯ ı-English dictionary. Lecture Notes in Computer Science, 3130:55–58, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Filippou:2004:HPA

[22] Dimitrios Filippou. Hyphenation patterns for ancient and modern Greek. Lecture Notes in Computer Science, 3130:59– 67, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Beesley:2004:TDA

[23] Kenneth R. Beesley. Typesetting the Deseret alphabet with L^{A}TEX and - . Lecture Notes in Computer Science, 3130:68–111, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goncalves:2004:FRM

[24] Luis Nobre Gonçalves. and a review of 3D packages. Lecture Notes in Computer Science, 3130: 112–124, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Padovani:2004:IEM

[25] Luca Padovani. Interactive editing of MathML markup using TEX syntax. Lecture Notes in Computer Science, 3130:

125–138, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cho:2004:TCL

[26] Jin-Hwan Cho and Haruhiko Okumura.

Typesetting CJK languages with Ω. Lecture Notes in Computer Science, 3130: 139–148, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eddahibi:2004:DAM

[27] Mustapha Eddahibi, Azzeddine Lazrek, and Khalid Sami. Dynamic Arabic mathematical fonts. Lecture Notes in Computer Science, 3130:149–157, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Eddahibi:2004:AMD

[28] Mostafa Banouni, Mohamed Elyaakoubi, and Azzeddine Lazrek. Arabic mathematical e-documents. Lecture Notes in Computer Science, 3130:158–168, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bzyl:2004:MXC

[29] Włodzimierz Bzyl and Tomasz Przechlewski. Migrating to XML: The case of the GUST Bulletin archive. Lecture Notes in Computer Science, 3130:169– 178, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holecek:2004:APG

[30] Jan Holeček and Petr Sojka. Animations in pdfTEX-generated PDF: A new method for directly embedding animation into PDF. Lecture Notes in Computer Science, 3130:179–191, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Laurens:2004:IIT

[31] Jérôme Laurens. iTEXmac: An integrated TEX environment for Mac OS X. Lecture Notes in Computer Science, 3130: 192–202, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hufflen:2004:MBL

[32] Jean-Michel Hufflen. MlBibTEX: Beyond L^{A}TEX. Lecture Notes in Computer Science, 3130:203–215, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Przechlewski:2004:MTR

[33] Tomasz Przechlewski. Managing TEX resources with XML topic maps. Lecture Notes in Computer Science, 3130:216– 228, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Schaefer:2004:SSC

[34] Frank-Rene Schaefer. ŞäferTEX: Source code esthetics for automated typesetting. Lecture Notes in Computer Science, 3130: 229–239, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Piska:2004:CTF

[35] Karel Píška. Creating Type 1 fonts from sources: Comparison of tools, techniques and results. Lecture Notes in Computer Science, 3130: 240–256, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Williams:2004:BGA

[36] George Williams. Beyond glyphs, advanced typographic features of fonts. Lecture Notes in Computer Science, 3130: 257–263, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biryukov:2004:MLA

[37] Alex Biryukov, Christophe De Cannière, and Michaël Quisquater. On multiple linear approximations. Lecture Notes in Computer Science, 3152:1–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Courtois:2004:FSB

[38] Nicolas T. Courtois. Feistel schemes and bi-linear cryptanalysis: (extended abstract). Lecture Notes in Computer Science, 3152:23–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boneh:2004:SGS

[39] Dan Boneh, Xavier Boyen, and Hovav Shacham. Short group signatures. Lecture Notes in Computer Science, 3152:41–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Camenisch:2004:SSA

[40] Jan Camenisch and Anna Lysyanskaya.

Signature schemes and anonymous credentials from bilinear maps. Lecture Notes in Computer Science, 3152:56–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Holenstein:2004:CCB

[41] Thomas Holenstein, Ueli Maurer, and Johan Sjödin. Complete classification of bilinear hard-core functions. Lecture Notes in Computer Science, 3152:73–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hsiao:2004:FCP

[42] Chun-Yuan Hsiao and Leonid Reyzin.

Finding collisions on a public road, or do secure hash functions need secret coins? Lecture Notes in Computer Science, 3152: 92–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Patarin:2004:SRF

[43] Jacques Patarin. Security of random Feistel schemes with 5 or more rounds. Lecture Notes in Computer Science, 3152: 106–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Okeya:2004:SBR

[44] Katsuyuki Okeya, Katja Schmidt-Samoa, Christian Spahn, and Tsuyoshi Takagi. Signed binary representations revisited. Lecture Notes in Computer Science, 3152: 123–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Scott:2004:CP

[45] Michael Scott and Paulo S. L. M. Barreto. Compressed pairings. Lecture Notes in Computer Science, 3152:140–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

vanDijk:2004:AOC

[46] Marten van Dijk and David Woodruff. Asymptotically optimal communication for torus-based cryptography. Lecture Notes in Computer Science, 3152:157–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gentry:2004:HCR

[47] Craig Gentry. How to compress rabin ciphertexts and signatures (and more). Lecture Notes in Computer Science, 3152: 179–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cheng:2004:BSD

[48] Qi Cheng. On the bounded sum-of-digits discrete logarithm problem in finite fields. Lecture Notes in Computer Science, 3152: 201–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

May:2004:CRS

[49] Alexander May. Computing the RSA secret key is deterministic polynomial time equivalent to factoring. Lecture Notes in Computer Science, 3152:213–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gennaro:2004:MTC

[50] Rosario Gennaro. Multi-trapdoor commitments and their applications to proofs of knowledge secure under concurrent man-in-the-middle attacks. Lecture Notes in Computer Science, 3152:220–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Crescenzo:2004:CRR

[51] Giovanni Di Crescenzo, Giuseppe Persiano, and Ivan Visconti. Constantround resettable zero knowledge with concurrent soundness in the bare publickey model. Lecture Notes in Computer Science, 3152:237–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Damgaard:2004:ZKP

[52] Ivan Damgård, Serge Fehr, and Louis Salvail. Zero-knowledge proofs and string commitments withstanding quantum attacks. Lecture Notes in Computer Science, 3152:254–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Bellare:2004:KEA

[53] Mihir Bellare and Adriana Palacio. The knowledge-of-exponent assumptions and 3-round zero-knowledge protocols. Lecture Notes in Computer Science, 3152: 273–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Biham:2004:NCS

[54] Eli Biham and Rafi Chen. Near-collisions of SHA-0. Lecture Notes in Computer Science, 3152:290–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Joux:2004:MIH

[55] Antoine Joux. Multicollisions in iterated hash functions. application to cascaded constructions. Lecture Notes in Computer Science, 3152:306–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Abe:2004:ASF

[56] Masayuki Abe and Serge Fehr. Adaptively secure Feldman VSS and applications to universally-composable threshold cryptography. Lecture Notes in Computer Science, 3152:317–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Katz:2004:ROS

[57] Jonathan Katz and Rafail Ostrovsky. Round-optimal secure two-party computation. Lecture Notes in Computer Science, 3152:335–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Landau:2004:SLE

[58] Susan Landau. Security, liberty, and electronic communications. Lecture Notes in Computer Science, 3152:355–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Molland:2004:ICA

[59] Håvard Molland and Tor Helleseth. An improved correlation attack against irregular clocked and filtered keystream generators. Lecture Notes in Computer Science, 3152:373–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Hawkes:2004:RVC

[60] Philip Hawkes and Gregory G. Rose.

Rewriting variables: The complexity of fast algebraic attacks on stream ciphers. Lecture Notes in Computer Science, 3152: 390–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Lu:2004:FCA

[61] Yi Lu and Serge Vaudenay. Faster correlation attack on Bluetooth keystream generator E0. Lecture Notes in Computer Science, 3152:407–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Kurosawa:2004:NPH

[62] Kaoru Kurosawa and Yvo Desmedt.

A new paradigm of hybrid encryption scheme. Lecture Notes in Computer Science, 3152:426–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boneh:2004:SIB

[63] Dan Boneh and Xavier Boyen. Secure identity based encryption without random oracles. Lecture Notes in Computer Science, 3152:443–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Moran:2004:NIT

[64] Tal Moran, Ronen Shaltiel, and Amnon Ta-Shma. Non-interactive timestamping in the bounded storage model. Lecture Notes in Computer Science, 3152:460–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Catalano:2004:IIP

[65] Dario Catalano, David Pointcheval, and Thomas Pornin. IPAKE: Isomorphisms for Password-Based Authenticated Key Exchange. Lecture Notes in Computer Science, 3152:477–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dodis:2004:REK

[66] Yevgeniy Dodis, Rosario Gennaro, Johan Håstad, Hugo Krawczyk, and Tal Rabin. Randomness extraction and key derivation using the CBC, Cascade and HMAC modes. Lecture Notes in Computer Science, 3152:494–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Goodrich:2004:ETB

[67] Michael T. Goodrich, Jonathan Z. Sun, and Roberto Tamassia. Efficient treebased revocation in groups of low-state devices. Lecture Notes in Computer Science, 3152:511–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dwork:2004:PPD

[68] Cynthia Dwork and Kobbi Nissim.

Privacy-preserving datamining on vertically partitioned databases. Lecture Notes in Computer Science, 3152:528–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Srinathan:2004:OPS

[69] K. Srinathan, Arvind Narayanan, and C. Pandu Rangan. Optimal perfectly secure message transmission. Lecture Notes in Computer Science, 3152:545–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fitzi:2004:PSB

[70] Matthias Fitzi, Stefan Wolf, and Jürg Wullschleger. Pseudo-signatures, broadcast, and multi-party computation from correlated randomness. Lecture Notes in Computer Science, 3152:562–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Sinderson:2004:MWS

[71] Elias Sinderson, Vish Magapu, and Ronald Mak. Middleware and Web services for the collaborative information portal of NASA’s Mars Exploration Rovers Mission. Lecture Notes in Computer Science, 3231:1–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:2004:CME

[72] Brian F. Cooper. A content model for evaluating peer-to-peer searching techniques. Lecture Notes in Computer Science, 3231:18–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cai:2004:FNL

[73] Hailong Cai and Jun Wang. Foreseer: A novel, locality-aware peer-to-peer system architecture for keyword searches. Lecture Notes in Computer Science, 3231:38–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Cooper:2004:GQI

[74] Brian F. Cooper. Guiding queries to information sources with InfoBeacons. Lecture Notes in Computer Science, 3231:59– ??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jelasity:2004:PSS

[75] Márk Jelasity, Rachid Guerraoui, Anne-Marie Kermarrec, and Maarten van Steen. The Peer Sampling Service: Experimental evaluation of unstructured Gossip-based implementations. Lecture Notes in Computer Science, 3231:79–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Maroti:2004:DFR

[76] Miklós Maróti. Directed flood-routing framework for wireless sensor networks. Lecture Notes in Computer Science, 3231: 99–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jin:2004:EPO

[77] Jingwen Jin and Klara Nahrstedt. On exploring performance optimizations in Web service composition. Lecture Notes in Computer Science, 3231:115–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Li:2004:ILM

[78] Baochun Li, Jiang Guo, and Mea Wang.

iOverlay: A lightweight middleware infrastructure for overlay application implementations. Lecture Notes in Computer Science, 3231:135–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Plattner:2004:GSR

[79] Christian Plattner and Gustavo Alonso.

Ganymed: Scalable replication for transactional Web applications. Lecture Notes in Computer Science, 3231:155–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Milan-Franco:2004:AMD

[80] Jesús M. Milan-Franco, Ricardo Jiménez-Peris, Marta Patiño-Martínez, and Bettina Kemme. Adaptive middleware for data replication. Lecture Notes in Computer Science, 3231:175–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Leff:2004:AES

[81] Avraham Leff and James T. Rayfield.

Alternative edge-server architectures for Enterprise JavaBeans applications. Lecture Notes in Computer Science, 3231: 195–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Nayate:2004:TID

[82] Amol Nayate, Mike Dahlin, and Arun Iyengar. Transparent information dissemination. Lecture Notes in Computer Science, 3231:212–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Wang:2004:OBP

[83] Jinling Wang, Beihong Jin, and Jing Li. An ontology-based publish/subscribe system. Lecture Notes in Computer Science, 3231:232–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gupta:2004:MCB

[84] Abhishek Gupta, Ozgur D. Sahin, Divyakant Agrawal, and Amr El Abbadi. Meghdoot: Content-based publish/ subscribe over P2P networks. Lecture Notes in Computer Science, 3231:254–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Zhao:2004:SPH

[85] Yuanyuan Zhao, Daniel Sturman, and Sumeer Bhola. Subscription propagation in highly-available publish/subscribe middleware. Lecture Notes in Computer Science, 3231:274–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tai:2004:CCW

[86] Stefan Tai, Rania Khalaf, and Thomas Mikalsen. Composition of coordinated Web services. Lecture Notes in Computer Science, 3231:294–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Naik:2004:ARS

[87] Vijay K. Naik and Swaminathan Sivasubramanian, Sr.iram Krishnan. Adaptive resource sharing in a Web services environment. Lecture Notes in Computer Science, 3231:311–??, 2004. CO-DEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ponnekanti:2004:IAI

[88] Shankar R. Ponnekanti and Armando Fox. Interoperability among independently evolving Web services. Lecture Notes in Computer Science, 3231:331–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Prasad:2004:SMT

[89] Sushil K. Prasad, Vijay Madisetti, Shamkant B. Navathe, Raj Sunderraman, Erdogan Dogdu, Anu Bourgeois, Michael Weeks, Bing Liu, Janaka Balasooriya, Arthi Hariharan, Wanxia Xie, Praveen Madiraju, Srilaxmi Malladi, Raghupathy Sivakumar, Alex Zelikovsky, Yanqing Zhang, Yi Pan, and Saied Belkasim. SyD: A middleware testbed for collaborative applications over small heterogeneous devices and data stores. Lecture Notes in Computer Science, 3231:352–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Roman:2004:DPR

[90] Manuel Roman and Nayeem Islam. Dynamically programmable and reconfigurable middleware services. Lecture Notes in Computer Science, 3231:372–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Ranganathan:2004:MML

[91] Anand Ranganathan, Jalal Al-Muhtadi, Shiva Chetan, Roy Campbell, and M. Dennis Mickunas. MiddleWhere: A middleware for location awareness in ubiquitous computing applications. Lecture Notes in Computer Science, 3231: 397–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Mohapatra:2004:GTA

[92] Shivajit Mohapatra and Nalini Venkatasubramanian. A game theoretic approach for power aware middleware. Lecture Notes in Computer Science, 3231:417–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Jordan:2004:EJT

[93] Mick Jordan, Grzegorz Czajkowski, Kirill Kouklinski, and Glenn Skinner. Extending a J2EE server with dynamic and flexible resource management. Lecture Notes in Computer Science, 3231:439–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Oberle:2004:DMS

[94] Daniel Oberle, Andreas Eberhart, Steffen Staab, and Raphael Volz. Developing and managing software components in an ontology-based application server. Lecture Notes in Computer Science, 3231: 459–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Tilevich:2004:PED

[95] Eli Tilevich and Yannis Smaragdakis. Portable and efficient distributed threads for Java. Lecture Notes in Computer Science, 3231:478–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Billig:2004:PIM

[96] Andreas Billig, Susanne Busse, Andreas Leicher, and other. Platform independent model transformation based on TRIPLE. Lecture Notes in Computer Science, 3231: 493–??, 2004. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).